OpenStudy (anonymous):

Greg has two porcelain spheres. The radii of the spheres are in a ratio of 3 : 4. What is the ratio between the volumes of the two spheres?

OpenStudy (anonymous):

Volume is proportional to the Radius to the 3rd power, therefore ratio of volumes will be ratios of Radii to the third power too. Rest is for you...

OpenStudy (anonymous):

can you show me an example

OpenStudy (anonymous):

V1=4/3*pi*R1^3 and V2=4/3*pi*R2^3, then V1/V2=(R1/R2)^3